OpenStudy (anonymous):

A hydrogen atom has an electron of mass m and a proton of mass M, M>>m. The reduced mass is defined to be: u= mM/(m+M) Given u as a series in m/M: The 1st order correction term is defined by including the linear term by no higher order terms. If m = M/1836, by what percentage does including the 1st order correction change the estimate u=m?

OpenStudy (unklerhaukus):

The geometric series for |x|<1 is \[\frac1{1-x}=\sum_{k=0}^\infty x^k=1+x+x^2+x^3+...\]

OpenStudy (unklerhaukus):

\[u= \frac{mM}{m+M}=\frac{m}{(m/M)+1}=\frac{m}{1-(-m/M)}\] \[m = M/1836\] is equivalent to \[m/M=1/1836\] so that |x| = |-m/M| < 1

OpenStudy (unklerhaukus):

\[u=\frac m{1-(-m/M)}\\=m\Big[1+(-m/M)+(-m/M)^2+...\Big]\\=\]